OpenStudy (judynaffziger):

Mike has $22 in his bank account and deposits $11.50 each week. At the same time, Tim has $218 in his account but is withdrawing $13 each week. When will the accounts have the same balance? How much money will each of them have after 12 weeks

OpenStudy (photon336):

you can easily represent this using a linear function do you know how to do this?

OpenStudy (judynaffziger):

No i dont.

OpenStudy (photon336):

alright so for mike he has originally 22 dollars in his account and deposits 11.50 each week We can represent each as a linear function y = m(x)+b \[Mike~m(s) = $22+11.50w \] \[Tim~~T(s) = $218-13w \]

OpenStudy (photon336):

finding how how much money each has after 12 weeks is easy: you just do this. if you notice w = the amount of weeks so you'd do this: \[22+11.50(12) = $$~\ mikes~account~after~12~weeks \] \[218-13(12) = $$~Tim's~account~after~12~weeks \] if you want to find out at what time they both have the same amount of money, you would set both equations equal to each-other: \[22+11.50w = 218-13w \]